Laurent series
1. Aşagıdaki fonksiyonu 1 <1: +11 < 3 bölgesinde Laurent SC 223-2)

the following function open the Laurent series in 1<|z+1|<3
Answer #1

To find the Laurent series off (2)=>7/2_,1<\2 +1<3 $(2)=7(2-2)-() Put z+1=t Hence 1<\t]<3 z=1-1 Put these values in (1),we ge

Now , < #| <3 ||>1,} <3 Hence | 1 1 1 1 1 1 4t-1 2 ft -124t -3 1 1 1 1 1 1 ਵੀਰ (+:++-ਕੈ(1+++ +++ ft)=(1++ਨੇ+-+ + )-ਵੈ1+:+ i